OpenStudy (anonymous):

A heated piece of metal cools according to the function c(x) = (.5)^(x − 7), where x is measured in hours. A device is added that aids in cooling according to the function h(x) = −x − 2. What will be the temperature of the metal after 2 hours?

OpenStudy (anonymous):

@PrincessHush

OpenStudy (anonymous):

@Deadly_Roses

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

@ganeshie8 @dan815

OpenStudy (campbell_st):

so all you do is substitute x = 2 into both equations so the temperature equation is t(x) = c(x) + h(x) so using x = 2 t(2) = c(2) + h(2) that should give an answer

OpenStudy (anonymous):

but where did the 2 come from?

OpenStudy (anonymous):

@Hero @ParthKohli

OpenStudy (anonymous):

@ikram002p

OpenStudy (campbell_st):

the 2 is gievn to you... you just need to substitute 2 into both equations to find the temp. \[Temp = (0.5)^{2 - 5} + (-2 - 2)\] and calculate the value. the simplified version is \[Temp = (0.5)^{-5} - 4\]